OpenStudy (anonymous):

lim as x->2 of sqrt(x^2-4)

sam (.sam.):

Just do it \[\lim_{x \rightarrow 0} \sqrt{x^2-4} \\ \\ =\lim_{x \rightarrow 0} \sqrt{4-4} \\ \\ =\lim_{x \rightarrow 0} \sqrt{0} \\ \\ =0\]